Money deposited in a certain bank doubles every 8 years. The bank compounds interest continuously. What annual interest rate does the bank offer? Type your answer as a percent (not a decimal), and round your answer to the nearest thousandth of a percent (3 decimal places).
Answer by Boreal(15235)

The rule of 72 suggests the answer will be close to 9 %.
P=Poexp(rt)
P/Po=2
2=exp(rt)
ln of both sides
ln2=rt
ln2/8=r
r=8.664 %